OpenStudy (anonymous):

which is an equation in slope-intercept form for the line passing through (5,1) and (-4,7)

OpenStudy (austinl):

Okay, first you need to get the slope between these two points, do you know how to do that?

OpenStudy (anonymous):

I have no idea :( math is my weak spot

OpenStudy (austinl):

\(\Large{m=slope=\frac{rise}{run}=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}}\)

OpenStudy (anonymous):

A. y=-2/3X-13/3 B. y=-2/3X+13/3 C. y=2/3X+13/3 D. y=2/3X-13/3

OpenStudy (accessdenied):

If dealing with the raw numbers isn't your cup of tea, maybe pictures would entertain you more? If you graph the points (5, 1) and (-4, 7), it might give you some insight on the equation that describes it. The characteristics will direct you to what kind of slope is happening and where the intercept is relative to the origin (0, 0). E.g.  You read this graph left-to-right, you can tell the y-value is decreasing, so its slope (tendency to change in y-values) has to be negative (decreasing values). The y-intercept is apparently positive because it is above the y-value 0. While I wouldn't necessarily recommend it for any type of these problems, the insight might help you understand the raw numbers a little better. :)

OpenStudy (accessdenied):

* Note that I did not graph the points from the problem in that graph, btw. Those were selected arbitrarily!